A charge q is distributed over two concentric hollow spheres of radii r and R(>r) such that the surface densities are equal. Find the potential at he common centre.

Grade:12

2 Answers

ananthupotti
7 years ago
charge density=q/4(3.14)r=Q/4(3.14)R.
ie, q=Q*r*r/R*R.
potential at center of shell=kq/r+kQ/R
kQ/R(r/R+1)
Q=charge density*4(3.14)R*R
on simpilfying,
V=charge density*(r+R)/epsilon
Rishi
6 years ago
As given:Surface charge density is equal meanssigma1=sigma2Let q` charge on rAnd Q on RThen q`/4πr^2 =Q/4πR^2q`/Q =r^2/R^2. ......(1) Now potential at common centre isV= kQ/R + kq`/rTaking kQ commonV= kQ(1/R+ q`/Qr)V= kQ(1/R +r/R^2)V= kQ{(r +R)/R^2}